using L hospital rule lim x=0 x-sinxcosx/x3

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
L = lim(x\rightarrow 0)\frac{x-sinx.cosx}{x^{3}}
L = lim(x\rightarrow 0)\frac{x-\frac{sin2x}{2}}{x^{3}}
L = lim(x\rightarrow 0)\frac{1-cos2x}{3x^{2}}
L = lim(x\rightarrow 0)\frac{2sin2x}{6x}
L = \frac{2}{3}lim(x\rightarrow 0)\frac{sin2x}{2x} = \frac{2}{3}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty